# On-Device Training of Fully Quantized Deep Neural Networks on Cortex-M Microcontrollers

### v1

@@ Abstract
On-device training of DNNs allows models to adapt and fine-tune to newly collected data or changing domains while deployed on microcontroller units (MCUs). However, DNN training is a resource-intensive task, making the implementation and execution of DNN training algorithms on MCUs challenging due to low processor speeds, constrained throughput, limited floating-point support, and memory constraints. In this work, we explore on-device training of DNNs for Cortex-M MCUs. We present a method that enables efficient training of DNNs completely in place on the MCU using fully quantized training (FQT) and dynamic partial gradient updates. We demonstrate the feasibility of our approach on multiple vision and time-series datasets and provide insights into the tradeoff between training accuracy, memory overhead, energy, and latency on real hardware.
@@ III-A Fully Quantized Training
$w_{i+1}=\frac{1}{s_{w_{i+1}}}\left[\left(w_{i}-z_{w_{i}}\right)s_{w_{i}}l\nabla w_{i}\right]+z_{w_{i+1}}$ (5)
adapted for quantization where $s_{w_{i}}$ and $z_{w_{i}}$ are the current weight tensor’s scale and zero point parameters and $s_{w_{i+1}}$ and $z_{w_{i+1}}$ are the updated weight tensor’s parameters. To update the scale and zero point parameters alongside the trainable weights, $s_{w_{i+1}}$, see Eq. (6), and $z_{w_{i+1}}$, see Eq. (7), are calculated based on the maximum and minimum values $f_{min}$ and $f_{max}$ observed in the intermediate results from $\left(w_{i}-z_{w_{i}}\right)s_{w_{i}}l\nabla w_{i}$ which are in floating-point space.
@@
To implement stochastic gradient descent (SGD), i.e., mini-batches, there are two viable approaches: (a) process all $b$ samples of a mini-batch at once during a single network pass by introducing an additional dimension to the input, output, and all intermediate tensors and then perform BP and weight updates immediately over the entire batch, or (b) accumulate $\nabla W_{n}$ in a buffer over $b$ successive training steps of individual samples. Both options increase the amount of memory required compared to performing gradient descent on each sample, i.e., batch size $b=1$. However, not using SGD and instead relying on per-sample updates can have a significant impact on the training performance of the DNN, as negative training effects such as catastrophic forgetting [20] become much more prevalent.
We have chosen version (b) for our SGD approach. The reason is that by introducing additional gradient buffers to accumulate $\nabla W_{n}$, the additional SRAM utilization to support SGD is independent of both the batch size $b$ and the full network depth $N$, and instead depends only on the number of layers actually trained in the DNN. Furthermore, to minimize the additional memory overhead introduced by SGD, we do not consider extensions such as (Nesterov-) momentum or adaptive learning rate algorithms such as Adam [21], as these techniques, while generally improving convergence and stability of training, come at the cost of significantly higher memory requirements due to the need to store multiple versions of the gradient buffer.
@@
$w_{i+1}=\frac{1}{s_{w_{i+1}}}\left[\left(w_{i}-z_{w_{i}}\right)s_{w_{i}}l\frac{\nabla w_{i}-\mu_{\nabla W}}{\sigma_{\nabla W}}\right]+z_{w_{i+1}}$ (8)
@@ IV-E Comparison with MCUNet-5FPS
MCUNet has more trainable parameters than MBedNet, especially in its last layers. This results in an overall higher memory utilization of MCUNet, see the left plot in Fig. 9. More specifically, since the last layers are trainable, this results in (a) larger feature maps to propagate between layers (blue segment of bars), and (b) more memory required to store larger gradient and weight buffers to enable training of the last layers (orange segment of bars). This trend can likewise be observed for latency, see the left plot in Fig. 9. Because MBedNet’s final layers are smaller, both their forward and backward passes can be executed significantly faster than MCUNet’s. Since the backward pass has to compute two partial derivatives for each layer, the difference is more pronounced than for the forward pass. Finally, when comparing accuracy, both architectures achieve a similar baseline of around 86%. When using our optimizer, MCUNet could be trained to an accuracy of 89.5% for cifar10, which is slightly higher than the 85.9% our optimizer achieved for MBedNet but which is in line with what SGD-M-QAS achieved on MCUNet (86.9%). Still, we argue that the significantly lower memory and latency cost of MBedNet compared to MCUNet makes it a more suitable architecture for efficient on-device training on MCUs. This allows MbedNet not only to run inference at around 5FPS (200 ms) like MCUNet, but also to retrain.

### v2

@@ Abstract
On-device training of DNNs allows models to adapt and fine-tune to newly collected data or changing domains while deployed on microcontroller units (MCUs). However, DNN training is a resource-intensive task, making the implementation and execution of DNN training algorithms on MCUs challenging due to low processor speeds, constrained throughput, limited floating-point support, and memory constraints. In this work, we explore on-device training DNNs for different sized Cortex-M MCUs (Cortex-M0+, Cortex-M4, and Cortex-M7). We present a method that enables efficient training of DNNs completely in place on the MCU using fully quantized training (FQT) and dynamic partial gradient updates. We demonstrate the feasibility of our approach on multiple vision and time-series datasets and provide insights into the tradeoff between training accuracy, memory overhead, energy, and latency on real hardware. The results show that compared to related work, our approach requires 34.8% less memory and has a 49.0% lower latency per training sample, with dynamic partial gradient updates allowing a speedup of up to 8.7 compared to fully updating all weights.
@@ III-A Fully Quantized Training
$w_{i+1}=\frac{1}{s_{w_{i+1}}}\left[\left(w_{i}-z_{w_{i}}\right)s_{w_{i}}-\ell\nabla w_{i}\right]+z_{w_{i+1}}$ (5)
adapted for quantization where $\ell$ is the learning rate, $s_{w_{i}}$ and $z_{w_{i}}$ are the current weight tensor’s scale and zero point parameters, and $s_{w_{i+1}}$ and $z_{w_{i+1}}$ are the updated weight tensor’s parameters. To update the scale and zero point parameters alongside the trainable weights, $s_{w_{i+1}}$, see Eq. (6), and $z_{w_{i+1}}$, see Eq. (7), are calculated based on the maximum and minimum values $f_{min}$ and $f_{max}$ observed in the intermediate results from $\left(w_{i}-z_{w_{i}}\right)s_{w_{i}}l\nabla w_{i}$ which are in floating-point space.
@@
To implement stochastic gradient descent (SGD), i.e., mini-batches, there are two viable approaches: (a) have all $b$ input samples of a mini-batch in RAM at the same time and process them all at once during a single network pass by introducing an additional dimension to the input, output, and all intermediate tensors, or (b) accumulate $\nabla W_{n}$ in a buffer over $b$ successive training steps of individual samples. Both options increase the amount of memory required compared to performing gradient descent on each sample, i.e., batch size $b=1$. However, not using SGD and instead relying on per-sample updates can have a significant impact on the training performance of the DNN, as negative training effects such as catastrophic forgetting [20] become much more prevalent.
We have chosen version (b) for our SGD approach. The reason is that by introducing additional gradient buffers to accumulate $\nabla W_{n}$, the additional SRAM utilization to support SGD is independent of both the batch size $b$ and the full network depth $N$, and instead depends only on the number of layers actually trained in the DNN. Most importantly, this approach does not require all the samples in a minibatch to be stored in RAM at the same time. Furthermore, to minimize the additional memory overhead introduced by SGD, we do not consider extensions such as (Nesterov-) momentum or adaptive learning rate algorithms such as Adam [21], as these techniques, while generally improving convergence and stability of training, come at the cost of significantly higher memory requirements due to the need to store multiple versions of the gradient buffer.
@@
$w_{i+1}=\frac{1}{s_{w_{i+1}}}\left[\left(w_{i}-z_{w_{i}}\right)s_{w_{i}}-\ell\frac{\nabla w_{i}-\mu_{\nabla W}}{\sigma_{\nabla W}}\right]+z_{w_{i+1}}$ (8)
@@ IV-E Comparison with MCUNet-5FPS
MCUNet has more trainable parameters than MBedNet, especially in its last layers. This results in (a) 34.8% less memory requirements and (b) 49.0% smaller latency of MBedNet compared to MCUNet, see Fig. 9. For memory, this results in (a) larger feature maps to propagate between layers (blue segment of bars), and (b) more memory required to store larger gradient and weight buffers to enable training of the last layers (orange segment of bars). This trend can likewise be observed for latency, see the left plot in Fig. 9. Because MBedNet’s final layers are smaller, both their forward and backward passes can be executed significantly faster than MCUNet’s. Since the backward pass has to compute two partial derivatives for each layer, the difference is more pronounced than for the forward pass. Finally, when comparing accuracy, both architectures achieve a similar baseline of around 86%. When using our optimizer, MCUNet could be trained to an accuracy of 89.5% for cifar10, which is slightly higher than the 85.9% our optimizer achieved for MBedNet but which is in line with what SGD-M-QAS achieved on MCUNet (86.9%). Still, we argue that the significantly lower memory and latency cost of MBedNet compared to MCUNet makes it a more suitable architecture for efficient on-device training on MCUs. This allows MbedNet not only to run inference at around 5FPS (200 ms) like MCUNet, but also to retrain.